OpenStudy (anonymous):

can someone help me on 2 sqrt 2 sin (theta) +2=0

OpenStudy (anonymous):

\[2 \sqrt{2 \sin \theta} + 2 = 0\] or \[2 \sqrt{2} \sin \theta + 2 = 0\]

OpenStudy (anonymous):

the second one

OpenStudy (anonymous):

How do you think you should start?

OpenStudy (anonymous):

I really dont know lol but i normally factor it out

OpenStudy (anonymous):

You could factor out a 2 first. But here you want to subtract the 2 first, then divide by 2√2

OpenStudy (anonymous):

so 2 sqrt 2 theta -2 = 0

OpenStudy (anonymous):

Subtract the 2 from both sides of the equation

OpenStudy (anonymous):

(2 √2) sin Θ = -2

OpenStudy (anonymous):

then factor?

OpenStudy (anonymous):

Then divide by 2√2

OpenStudy (anonymous):

You don't have to factor this equation at all

OpenStudy (anonymous):

sin theta = -1

OpenStudy (anonymous):

Sorry I got kicked out & now experiencing a lot of lag...

OpenStudy (anonymous):